OpenStudy (anonymous):

Find the derivatuve f(x)=-7/x at x=-3

OpenStudy (anonymous):

\[\frac{\partial }{\partial x}\left(-\frac{7}{x}\right)=\frac{7}{x^2} \]At x = -3, 7/x^2 = 7/9

OpenStudy (anonymous):

Oh okay I get it now thanks!

OpenStudy (anonymous):

Your welcome. The notation is for taking the partial derivative. x is the only variable so the partial derivative is the normal derivative.